# Methods Paper Version 19

### v1

@@ Abstract
We apply the Hulthén–Kohn method suggested by V. D. Efros [Phys. Rev. C 99, 034620 (2019)] for calculating various observables in the continuum and discrete spectrum using two-body interactions in single- and coupled-channel systems. This method is promising for many-body applications and ab initio description of nuclear reactions. We explore the convergence of phase shifts and wave functions as well as the location of $S$-matrix poles which enables obtaining both resonance and bound state parameters. We find that adopting wave functions from approximate bound-state solutions for the short-range components of basis wave functions leads to good convergence.
@@ I Introduction
Over the last 25 years, various ab initio (from first principles) methods have been developed and extensively applied with the aim of describing the structure of atomic nuclei, including the Green’s function Monte Carlo method GFMC0 ; GFMC1 ; GFMC2 ; GFMC3 ; GFMC4 ; GFMC5 , the coupled cluster method CoesterOne ; CoesterTwo ; Kummel ; Wloh ; Hagen ; Hagen2022 ; ZHSun2023 , and the no-core shell model (NCSM) NCSMPRL ; NCSMProg . In principle, one can choose any basis within the Hilbert space to construct the eigenfunctions. For practical reasons, the harmonic oscillator basis is conventionally used in nuclear structure theory as it possesses several convenient features for applications to realistic many-body problems. The utility is such that the many-body problem is reduced within the M-scheme (in which all basis states have a fixed value of total angular momentum projection but not a fixed total angular momentum) to a sparse matrix eigenvalue problem that is solvable with leadership class supercomputers MFDnMaris2010 ; Euro-Par ; Aktulga ; NCSM-NTSE2014 ; Shao .
@@
In light of these needs, several ab initio methods have been applied to elastic scattering problems. For example, the Green’s function Monte Carlo has been used to obtain features of the $J^{\pi}=\frac{3}{2}^{-}$ and $\frac{1}{2}^{-}$ low-lying resonant states in $n\alpha$ scattering Nollett . Properties of nuclear resonances and reactions with light nuclei can also be studied within the no-core Gamow shell model Papadimitriou . In particular, this approach was successfully applied recently to the study of the tetraneutron and trineutron JLi2019 . Another successful approach is the resonating group method (RGM) in combination with the NCSM QuagOne ; QuagTwo ; Navratil2010 , which has been used to study nuclear scattering in systems as large as $A=17$ Navratil2010 , and its more extended version, the NCSM with continuum BaroniTwo ; BaroniOne ; Navr-scr ; NCSMCHebborn2022 ; NCSMCGysbers2023 .
@@
An alternative approach to the studies of the continuum spectrum is the harmonic oscillator representation of scattering equations (HORSE) BangHORSE ; Zaitsev1998 . The HORSE method allows one to calculate the continuum wave function and the scattering phase shifts $\delta_{l}(E)$. It has already been successfully applied to the calculations of $np$ scattering and the $M1$ transition cross section of the reaction $p(n,\gamma)d$ DuWeijieHORSE  as well as to calculations of continuum nuclear spectra within various cluster models with phenomenological interactions in Refs. monopoleC12 ; Lurie1 ; Lurie2 ; Lurie3 ; Lurie4 and within the RGM 6He6He ; VVasil ; Lashko . For many-body ab initio applications, however, the HORSE method is impractical due to the prohibitive computational cost. To get around this, a simplified HORSE version, the single-state HORSE (SS-HORSE) method has been developed AMShirokovHORSE2016 ; BlokhintsevSSHorseC ; AMShirokovHORSE2018 , which makes it possible to extend the NCSM to calculations of phase shifts. Moreover, the SS-HORSE method can also provide resonance energies and widths via $S$-matrix pole location. The SS-HORSE method in combination with the NCSM was used to describe resonances in various light nuclei, in particular, in 5He, 5Li AMShirokovHORSE2016 ; AMShirokovHORSE2018 , 7He PPN ; Igor2022 , 9Li 9Li , to propose the first description of the tetraneutron resonance AMShirokov2016 ; 4nAIP , and to predict resonant states in the trineutron 3narXiv . The price paid for the SS-HORSE simplification is the loss of information on the continuum wave function. Without the wave functions, one is not able to perform calculations of important physical observables such as, e. g., electromagnetic dissociation cross sections.
@@
The modified Hulthén–Kohn method extending the NCSM to the continuum spectrum was suggested by V. D. Efros VDEfros to address the shortcomings of the HORSE and SS-HORSE approaches in a computationally feasible manner while retaining the predictive power of the ab initio nature of the NCSM. Moreover, unlike the SS-HORSE method, the Efros method easily extends to multichannel problems. To date, this approach has been tested using simple model problems only in Ref. VDEfros .
The aim of this work is to explore in detail the possibilities of the Efros method VDEfros in solving scattering problems with the goal of utilizing it in future many-body applications. We employ this improved formalism to calculate and analyze phase shifts and continuum wave function convergence for both single and coupled channels using the Daejeon16 $NN$ interaction Daejeon16  and model two-body problems. We demonstrate the utilization of the Efros method for locating $S$-matrix poles to find resonance energies $E_{r}$ and widths $\Gamma$ as well as to improve variational results for the bound state energies $E_{b}$. We combine the ideas of the Efros method with some achievements of the HORSE formalism BangHORSE ; Zaitsev1998 which is a particular version of the $J$-matrix formalism Yamani  in quantum scattering theory. The $J$-matrix utilizes not only the oscillator basis expansions but also those in terms of the Laguerre functions which make it possible to include naturally the Coulomb interactions in the calculations. For many-body nuclear applications, however, the basis of Laguerre functions is not as convenient as the harmonic oscillator basis and we therefore focus here on the HORSE version of the $J$-matrix. As a result, we obtain both continuum and bound state wave functions as infinite series of oscillator functions. We discuss here only the scattering of uncharged particles; various extensions of the HORSE formalism accounting for the Coulomb distortion of wave function asymptotics can be found in Refs. Okhrimenko ; BangHORSE .
@@
In Section II, we introduce the formalism of the Efros method. We discuss applications of this method to single- and coupled-channel $np$ scattering in Sections III and IV, respectively. Calculations of $S$-matrix poles are examined in Section V. We conclude in Section VI.
@@ II Formalism
We start with presenting the Efros method VDEfros  in the case of a two-body problem with $w$ open channels and some modifications and extensions that we suppose will be useful for many-body applications.
@@ II.1 Ansatz wave function
Here $\vec{r}$ is the coordinate of relative separation of colliding fragments, the channel multi-indices $i$ and $j$ include all quantum numbers distinguishing the channels and those common for all coupled channels like the orbital momentum of relative motion of colliding particles $l_{j}$ in the channel $j$, their total spin $S_{j}$, and the total angular momentum $J$ and its projection $M_{J}$. $\beta_{q}(\vec{r})$ are $v$ linearly-independent short-range functions (SRFs) vanishing as ${r\rightarrow\infty}$ which are introduced to describe the wave function at short distances. The set of $v$ energy-dependent coefficients $b_{q}^{i}(E)$, $q=0,...\>,v-1$, together with $w$ $S$-matrix elements $\mathbb{S}_{ji}(E)$, $j=1,...\>,w$, comprise a set of $N=v+w$ unknowns in each given entrance channel $i$ that will be determined within this approach.
For nuclear physics applications, it is natural and convenient to use the expansions in terms of harmonic oscillator functions; therefore we set SRFs $\beta_{q}(\vec{r})$ to be equal to finite sums of these functions,
@@
$\phi_{nj}(\vec{r})\equiv\phi_{nl_{j}}\!(\vec{r})=R_{nl_{j}}\!(r)\,\mathscr{Y}_{JM_{J}}^{l_{j}S_{j}}(\Omega_{\hat{r}}),$ (3)
where $\mathscr{Y}_{JM_{J}}^{l_{j}S_{j}}(\Omega_{\hat{r}})$ are generalized spherical harmonics which describe the spin-angular degrees of freedom in the channel $j$,
@@
$(lm_{l}Sm_{S}|JM_{J})$ are real-valued Clebsch–Gordan coefficients Varsha , $\chi_{m_{S}}^{S}$ is the spinor, and $Y_{lm_{l}}(\Omega_{\hat{r}})$ is the spherical harmonic function Varsha ;
@@
The irregularity at the origin of the spherical Neumann $n_{l}(z)$ and Hankel $h^{(1,2)}_{l}(z)$ functions Arfken  should be regularized to avoid divergences of integrals calculated with them.
@@
Therefore, contrary to Efros VDEfros  who uses regulator functions for the regularization of the functions $n_{l}(z)$ and $h^{(1,2)}_{l}(z)$, we utilize here the functions developed within the HORSE formalism BangHORSE ; Zaitsev1998 defined through their infinite expansions in a series of harmonic oscillator functions both for the regularized functions $\tilde{n}_{l}(kr)$ and $\tilde{h}^{\pm}_{l}(kr)$ and for the regular spherical Bessel function $j_{l}(kr)$:
@@
×$\phantom{\Gamma(-l)}\times{{}_{1}F_{1}}\Big{(}\!-n-l-\frac{1}{2},-l+\frac{1}{2};k^{2}b^{2}\Big{)},$ (17)
$C^{\pm}_{nl}(k)=$ $\>C_{nl}(k)\pm iS_{nl}(k),$ (18)
@@
${u}_{i}(E;\vec{r})=j_{l_{i}}(kr)\,\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})\\ +\sum_{j=1}^{w}K_{ji}(E)\,\tilde{n}_{l_{j}}(kr)\,\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})+\sum_{q=0}^{v-1}{b}_{q}^{i}(E)\,\beta_{q}(\vec{r}),$ (19)
@@ II.2 Formal solutions
[…]
@@
Efros suggests VDEfros to use the same $v$ SRFs $\bar{\beta}_{q}(\vec{r})=\beta_{q}(\vec{r})$, $q=0,...\,,v-1,$ and we shall mostly follow this suggestion in what follows. In this case, $\bar{a}_{qn;i}^{j}=a_{qn}^{j}$ for $q=0,...\,,{v-1}$. For $q=v,...\,,N$, $\bar{\beta}_{q}(\vec{r})$ may be arbitrarily chosen with the only constraint being that the functions $\bar{\beta}_{q}(\vec{r}),q=0,...,N$ should be linearly-independent. One of the reasonable choices, for example, may be some oscillator functions as we shall see below.
However, generally, the set of SRFs $\{\bar{\beta}_{q}\}$ can differ completely from the set of linearly-independent SRFs $\{{\beta}_{q}\}$; it is only important to have a set of $N$ linearly-independent SRFs $\bar{\beta}_{q}(\vec{r})$ which may be non-orthogonal. That means that, generally, the coefficients $\bar{a}_{qn;i}^{j}$ are distinct from $a_{qn}^{j}$ in Eq. (2). Moreover, $\bar{M}_{j}$, the finite truncations of oscillator series in each channel $j$, may differ from the truncations $M_{j}$ in Eq. (2).
@@
$\bar{A}_{qq^{\prime}}={\langle{\bar{\beta}_{q}}|H-E|\eta^{+}_{{j}}\rangle},\quad$$\begin{cases}{q=0,...\,,N-1},\\ q^{\prime}=v,...\,,N-1,\\ {j=q^{\prime}-v+1,}\end{cases}$ (26b)
@@
By solving Eq. (27) we can obtain all $wN$ unknowns entering Eqs. (1) with different entrance channels $i=1,...\,,w$. In the single-channel case, the above equations provide a complete solution of the problem. However, not all the $Nw$ unknowns are independent in the multichannel case when $w>1$. In particular, the $S$ matrix should be symmetric, i. e., $\mathbb{S}_{ij}(E)=\mathbb{S}_{ji}(E)$. So, the $S$ matrix has $w^{2}$ matrix elements but only ${\frac{1}{2}w(w+1)}$ independent matrix elements. Therefore, Eq. (27) is over-complete, providing ${w^{2}-\frac{1}{2}w(w+1)}=\frac{1}{2}w(w-1)$ unneeded extra relations between unknowns, and the number of independent unknowns is ${\mathbb{\color[rgb]{0,0,0}N}}{=Nw-\frac{1}{2}w(w-1)}$. We shall discuss this situation below using an example of a two-channel system. We shall show that in some cases the over-complete Eq. (27) provides a correct solution with $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)$. However, in cases of a reduced set of SRFs $\bar{\beta}_{q}$ most interesting for many-body applications, solving Eq. (27) results in $\mathbb{S}_{12}(E)\neq\mathbb{S}_{21}(E)$.
@@
Instead of $wN$ equations (22) we need now to formulate only $\mathbb{\color[rgb]{0,0,0}N}$ independent equations of the same type determining the $\mathbb{\color[rgb]{0,0,0}N}$ independent unknowns. We can do this by reducing the sets of bra SRFs $\langle{\bar{\beta}_{q}}|$ in Eq. (22) in some entrance channels $i$. We denote these reduced sets as $\{\langle\bar{\beta}_{q}^{i}|\}$. We include in the set $\{\langle\bar{\beta}_{q}^{1}|\}$ in the first entrance channel ($i=1$) all $N$ SRFs $\langle{\bar{\beta}_{q}}|$, $\{\langle\bar{\beta}_{q}^{1}|\}\equiv\{\langle{\bar{\beta}_{q}}|\}$. We note here that the channel ordering is arbitrary and we can always choose a particular numbering of channels which is optimal from the point of view of providing the best convergence. For the second entrance channel with $i=2$, we use $N-1$ independent SRFs $\langle\bar{\beta}_{q}^{2}|$ which may be obtained from the set $\{\langle{\bar{\beta}_{q}}|\}$ by excluding from it one of the SRFs $\langle{\bar{\beta}_{q}}|$. We can also define some of the SRFs $\langle\bar{\beta}_{q}^{2}|$ as linear combinations of the SRFs $\langle{\bar{\beta}_{q}}|$: the set $\{\langle\bar{\beta}_{q}^{2}|\}$ should include $N-1$ independent SRFs which can be non-orthogonal. We note that any superposition of $v$ linearly-independent SRFs $\beta_{q}(\vec{r})$, $q=0,...\,,v-1$, entering Eq. (1) should produce exactly the same results while superpositions involving additional $w$ SRFs ${\bar{\beta}_{q}(\vec{r})}$, $q=v,...\,,v+w$, involved in Eq. (22) may affect the convergence. Next, we obtain in the same manner the set $\{\langle\bar{\beta}_{q}^{3}|\}$ including $N-2$ independent SRFs $\langle\bar{\beta}_{q}^{3}|$, etc. We note that the set of ket SRFs $\{|\beta_{q^{\prime}}\rangle\}$ is the same in all entrance channels.
@@
$B^{1}_{q}=-\langle\bar{\beta}^{1}_{q}|H-E|\eta^{-}_{{i}}\rangle,\;\qquad$ $q=0,...\,,N-1,$ (34c)
@@
is an $\mathbb{\color[rgb]{0,0,0}N}\times\mathbb{\color[rgb]{0,0,0}N}$ matrix comprising $(N-i+1)\times(N-i+1)$ matrices $A^{i}$ as diagonal blocks,

### v2

@@ Abstract
We adapt the Hulthén–Kohn method suggested by Efros [Phys. Rev. C 99, 034620 (2019)] for calculating various observables in the continuum and discrete spectrum using two-body interactions in single- and coupled-channel systems. We explore the convergence of phase shifts and wave functions as well as the location of $S$-matrix poles which enables obtaining both resonance and bound state parameters. We find that employing a harmonic oscillator basis, together with an interaction smoothing scheme introduced by Gyarmati et al. [Nucl. Phys. A 326, 119 (1979)], and adopting approximate bound-state solutions for the short-range components of basis wave functions lead to good convergence even with restricted oscillator quanta accessible for modern no-core shell model (NCSM) codes. The adapted Efros method will facilitate ab initio many-body nuclear structure applications.
@@ I Introduction
Over the past 25 years, various ab initio (from first principles) methods have been developed and extensively applied with the aim of describing the structure of atomic nuclei, including the Green’s function Monte Carlo method GFMC0 ; GFMC1 ; GFMC2 ; GFMC3 ; GFMC4 , the coupled cluster method CoesterOne ; CoesterTwo ; Kummel ; Wloh ; Hagen ; Hagen2022 ; ZHSun2023 , and the no-core shell model (NCSM) NCSMPRL ; NCSMProg . In principle, one can choose any basis within the Hilbert space to construct the eigenfunctions. For practical reasons, the harmonic oscillator basis is conventionally used in nuclear structure theory as it possesses several convenient features for applications to realistic many-body problems. The utility is such that the many-body problem is reduced within the M-scheme (in which all basis states have a fixed value of total angular momentum projection but not a fixed total angular momentum) to a sparse matrix eigenvalue problem that is solvable with leadership class supercomputers MFDnMaris2010 ; Euro-Par ; Aktulga ; NCSM-NTSE2014 ; Shao .
@@
In light of these needs, several ab initio methods have been applied to elastic scattering problems. For example, the Green’s function Monte Carlo has been used to obtain features of the $J^{\pi}=\frac{3}{2}^{-}$ and $\frac{1}{2}^{-}$ low-lying resonant states in $n\alpha$ scattering Nollett . The quantum Monte Carlo method has been used to study $n-{{}^{3}\textrm{H}}$ scattering FloresOne and $n\alpha$ scattering FloresTwo . Properties of nuclear resonances and reactions with light nuclei can also be studied within the no-core Gamow shell model Papadimitriou . In particular, this approach was successfully applied recently to the study of the tetraneutron and trineutron JLi2019 . Another successful approach is the resonating group method (RGM) in combination with the NCSM QuagOne ; QuagTwo ; Navratil2010 , which has been used to study nuclear scattering in systems as large as $A=17$ Navratil2010 , and its more extended version, the NCSM with continuum BaroniTwo ; BaroniOne ; Navr-scr ; NCSMCHebborn2022 ; NCSMCGysbers2023 .
@@
An alternative approach to the studies of the continuum spectrum is the harmonic oscillator representation of scattering equations (HORSE) Zaitsev1998 ; BangHORSE . The HORSE method allows one to calculate the continuum wave function and the scattering phase shifts $\delta_{l}(E)$. It has already been successfully applied to the calculations of $np$ scattering and the $M1$ transition cross section of the reaction $p(n,\gamma)d$ DuWeijieHORSE  as well as to calculations of continuum nuclear spectra within various cluster models with phenomenological interactions in Refs. monopoleC12 ; Lurie1 ; Lurie2 ; Lurie3 ; Lurie4 and within the RGM 6He6He ; VVasil ; Lashko . For many-body ab initio applications, however, the HORSE method is impractical due to the prohibitive computational cost. To get around this, a simplified HORSE version, the single-state HORSE (SS-HORSE) method has been developed AMShirokovHORSE2016 ; BlokhintsevSSHorseC ; AMShirokovHORSE2018 , which makes it possible to extend the NCSM to calculations of phase shifts. Moreover, the SS-HORSE method can also provide resonance energies and widths via $S$-matrix pole location. The SS-HORSE method in combination with the NCSM was used to describe resonances in various light nuclei, in particular, in 5He, 5Li AMShirokovHORSE2016 ; AMShirokovHORSE2018 , 7He PPN ; Igor2022 , 9Li 9Li , to propose the first description of the tetraneutron resonance AMShirokov2016 ; 4nAIP , and to predict resonant states in the trineutron 3narXiv . The price paid for the SS-HORSE simplification is the loss of information on the continuum wave function. Without the wave functions, one is not able to perform calculations of important physical observables such as, e.g., electromagnetic dissociation cross sections.
@@
To address the shortcomings of the HORSE and SS-HORSE approaches in a computationally feasible manner, the Hulthén–Kohn method originally introduced in Refs. Hulthen1948 ; Kohn1948 ; Moses1953 ; Moses was modified by Efros VDEfros in a way that makes it convenient for extending the NCSM to the continuum spectrum. Moreover, unlike the SS-HORSE method, the Efros method easily extends to multichannel problems. To date, this approach has been tested using simple model problems only in Ref. VDEfros . Recently, Kohn’s variational method has also been used to solve the problem of scattering of nucleons by 3H or 3He within the hyperspherical harmonics approach Viviani2020 and has also been successfully utilized to solve fission reactions HaginoBertsch2024 .
In this work, we adapt the Efros method VDEfros for solving scattering problems.  We introduce additional features that improve the method with the goal of utilizing it in future many-body applications.  The adapted Efros method incorporates achievements of the HORSE formalism Zaitsev1998 ; BangHORSE which is a particular version of the $J$-matrix formalism Yamani  in quantum scattering theory that utilizes the oscillator basis. We use the HORSE formalism because the harmonic oscillator basis is natural and convenient for nuclear physics applications. Working in the harmonic oscillator basis, we construct basis wave functions by using different choices of linearly independent superpositions of oscillator functions as short-range components. For many-body calculations, we may choose as short-range components low-lying NCSM eigenfunctions that are obtainable from modern NCSM codes. We also discuss strategies for implementing our approach to solve ab initio many-body problems with restricted oscillator quanta accessible for modern NCSM codes. To further improve convergence, we utilize an interaction smoothing scheme introduced in Refs. HungOne ; HungTwo .
We employ this improved formalism to calculate and analyze phase shifts and continuum wave functions for both single and coupled channels using the Daejeon16 $NN$ interaction Daejeon16  and model two-body problems. We demonstrate the utilization of the adapted Efros method for locating $S$-matrix poles to find resonance energies $E_{r}$ and widths $\Gamma$ as well as to improve variational results for the bound state energies $E_{b}$. With these calculations, we demonstrate that the adapted Efros method leads to accurate results even with restricted oscillator quanta. We envision that our improvements to the method will facilitate ab initio many-body calculations of nuclear structure and reactions.
@@
In Section II, we introduce the formalism of the adapted Efros method. We discuss applications of this method to single- and coupled-channel $np$ scattering in Sections III and IV, respectively. Calculations of $S$-matrix poles are examined in Section V. We conclude in Section VI. Additional results are presented in Supplemental Material (SM) Supplemental .
@@ II Formalism
We start with presenting the adapted Efros method VDEfros  in the case of a two-body problem with $w$ open channels and some modifications and extensions that we suppose will be useful for many-body applications.
@@ II.1 Ansatz wave function
Here $\vec{r}$ is the coordinate of relative separation of colliding fragments, the channel multi-indices $i$ and $j$ include all quantum numbers distinguishing the channels and those common for all coupled channels like the orbital momentum of relative motion of colliding particles $l_{j}$ in the channel $j$, their total spin $S_{j}$, and the total angular momentum $J$ and its projection $M_{J}$. $\beta_{q}(\vec{r})$ are $v$ linearly independent short-range functions (SRFs) vanishing as $r\rightarrow\infty$} which are introduced to describe the wave function at short distances. The set of $v$ energy-dependent coefficients $b_{q}^{i}(E)$, q=0,,v1q=0,...\>,v-1, together with $w$ $S$-matrix elements $\mathbb{S}_{ji}(E)$, j=1,,wj=1,...\>,w, comprise a set of $N=v+w$ unknowns in each given entrance channel $i$ that should be determined within this approach.
For nuclear physics applications, it is natural and convenient to use the expansions in terms of harmonic oscillator functions. The harmonic oscillator basis has been used in many ab initio nuclear physics applications, for example, in the NCSM and coupled cluster applications mentioned above as well as to describe clustering from first principles in $^{8}\textrm{Be}$, $^{10}\textrm{Be}$, and $^{12}\textrm{C}$ KravvarisVolya2017 . We set SRFs $\beta_{q}(\vec{r})$ to be equal to finite sums of harmonic oscillator functions,
@@
[…]
@@
[…]
@@
The singularity at the origin of the spherical Neumann $n_{l}(z)$ and Hankel $h^{(1,2)}_{l}(z)$ functions Arfken is regularized in the functions $\tilde{n}_{l}(z)$ and $\tilde{h}^{\pm}_{l}(z)$ that makes it possible to avoid divergences of integrals calculated with them.
@@
Therefore, contrary to Efros VDEfros  who uses regulator functions for the regularization of the functions $n_{l}(z)$ and $h^{(1,2)}_{l}(z)$, we utilize here the functions developed within the HORSE formalism Zaitsev1998 ; BangHORSE defined through their infinite expansions in a series of harmonic oscillator functions both for the regularized functions $\tilde{n}_{l}(kr)$ and $\tilde{h}^{\pm}_{l}(kr)$ and for the regular spherical Bessel function $j_{l}(kr)$:
@@
$\phantom{\Gamma(-l)}\times{{}_{1}F_{1}}\Big{(}\!-n-l-\frac{1}{2},-l+\frac{1}{2};k^{2}b^{2}\Big{)},$ (17)
$\color[rgb]{0,0,0}\color[rgb]{0,0,0}C^{\pm}_{nl}(k)=$ $\color[rgb]{0,0,0}\>C_{nl}(k)\pm iS_{nl}(k),$ (18)
@@
${u}_{i}(E;\vec{r})=j_{l_{i}}(kr)\,\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})\\ +\sum_{j=1}^{w}K_{ji}(E)\,\tilde{n}_{l_{j}}(kr)\,\mathscr{Y}_{JM_{J}}^{l_{j}S_{j}}(\Omega_{\hat{r}})+\sum_{q=0}^{v-1}{b}_{q}^{i}(E)\,\beta_{q}(\vec{r}),$ (19)
@@
Our method can incorporate Coulomb interactions in a straightforward manner by following Ref. BangHORSE . In particular, for the scattering of charged particles, we introduce an auxiliary interaction operator $V^{sh}$. The Coulomb interaction is included in $V^{sh}$ and is formally cut at a distance $r_{c}$ that is greater than the range of the nuclear interaction. Using the auxiliary interaction, we obtain the scattering wave function $u_{i}^{sh}(E;\vec{r})$. The solutions corresponding to $V^{sh}$ are then utilized to calculate the Coulomb-distorted scattering wave function $u_{i}^{Coul}(E;\vec{r})$ that is obtained via trivial recalculations. Additional details are presented in Ref. BangHORSE and alternative approaches can be found in Refs. Okhrimenko and Yanikov2025 .
@@ II.2 Formal solutions
$\langle\bar{\beta}_{q}|H-E|u_{i}(E;\vec{r})\rangle=0,\hskip 2.84526ptq=0,...\,,N-1;\hskip 2.84526pti=1,...\,,w.$ (22)
Here, $\bar{\beta}_{q}=\bar{\beta}_{q}(\vec{r})$ are SRFs. We note that the number of unknowns $N=v+w$ in Eq. (1) is larger than $v$, the number of SRFs $\beta_{q}(\vec{r})$ involved in this equation. Hence we need to use in Eq. (22ww more SRFs $\bar{\beta}_{q}(\vec{r})$ as compared with the SRFs $\beta_{q}(\vec{r})$ used to describe the short-range physics in Eq. (1). As in Eq. (2), we can expand $\bar{\beta}_{q}^{i}(\vec{r})$ in a finite series of oscillator functions,
@@
Efros suggests VDEfros to use the same $v$ SRFs $\bar{\beta}_{q}(\vec{r})=\beta_{q}(\vec{r})$, q=0,,v1,q=0,...\,,v-1, and we shall mostly follow this suggestion in what follows. In this case, $\bar{a}_{qn;i}^{j}=a_{qn}^{j}$ for q=0,,v1q=0,...\,,{v-1}. For q=v,,Nq=v,...\,,N, $\bar{\beta}_{q}(\vec{r})$ may be arbitrarily chosen with the only constraint being that the functions $\bar{\beta}_{q}(\vec{r}),q=0,...,N$ should be linearly independent. One of the reasonable choices, for example, may be some oscillator functions as we shall see below.
However, generally, the set of SRFs $\{\bar{\beta}_{q}\}$ can differ completely from the set of linearly independent SRFs $\{{\beta}_{q}\}$; it is only important to have a set of $N$ linearly independent SRFs $\bar{\beta}_{q}(\vec{r})$ which may be nonorthogonal. That means that, generally, the coefficients $\bar{a}_{qn;i}^{j}$ are distinct from $a_{qn}^{j}$ in Eq. (2). Moreover, $\bar{M}_{j}$, the finite truncations of oscillator series in each channel $j$, may differ from the truncations $M_{j}$ in Eq. (2).
@@
$\bar{A}_{qq^{\prime}}={\langle{\bar{\beta}_{q}}|H-E|\eta^{+}_{{j}}\rangle}$$\quad$ $\begin{cases}{q=0,...\,,N-1},\\ q^{\prime}=v,...\,,N-1,\\ {j=q^{\prime}-v+1,}\end{cases}$ (26b)
@@
By solving Eq. (27) we can obtain all $wN$ unknowns entering Eqs. (1) with different entrance channels i=1,,wi=1,...\,,w. In the single-channel case, the above equations provide a complete solution of the problem. However, not all the $Nw$ unknowns are independent in the multichannel case when $w>1$. In particular, the $S$ matrix should be symmetric, i. e., $\mathbb{S}_{ij}(E)=\mathbb{S}_{ji}(E)$. So, the $S$ matrix has $w^{2}$ matrix elements but only $\frac{1}{2}w(w+1)$} independent matrix elements. Therefore, Eq. (27) is over-complete, providing ${w^{2}-\frac{1}{2}w(w+1)}=\frac{1}{2}w(w-1)$ unneeded extra relations between unknowns, and the number of independent unknowns is ${\mathbb{\color[rgb]{0,0,0}N}}{=Nw-\frac{1}{2}w(w-1)}$. In Ref. (Supplemental, , Section III), we discuss this situation using an example of a two-channel system. We show that in some cases the over-complete Eq. (27) provides a correct solution with $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)$. However, in cases of a reduced set of SRFs $\bar{\beta}_{q}$ most interesting for many-body applications, solving Eq. (27) results in $\mathbb{S}_{12}(E)\neq\mathbb{S}_{21}(E)$.
@@
Instead of $wN$ equations (22) we need now to formulate only $\mathbb{\color[rgb]{0,0,0}N}$ independent equations of the same type determining the $\mathbb{\color[rgb]{0,0,0}N}$ independent unknowns. We can do this by reducing the sets of bra SRFs $\langle{\bar{\beta}_{q}}|$ in Eq. (22) in some entrance channels $i$. We denote these reduced sets as $\{\langle\bar{\beta}_{q}^{i}|\}$. We include in the set $\{\langle\bar{\beta}_{q}^{1}|\}$ in the first entrance channel ($i=1$) all $N$ SRFs $\langle{\bar{\beta}_{q}}|$, $\{\langle\bar{\beta}_{q}^{1}|\}\equiv\{\langle{\bar{\beta}_{q}}|\}$. We note here that the channel ordering is arbitrary and we can always choose a particular numbering of channels which is optimal from the point of view of providing the best convergence. For the second entrance channel with $i=2$, we use N1N-1 independent SRFs $\langle\bar{\beta}_{q}^{2}|$ which may be obtained from the set $\{\langle{\bar{\beta}_{q}}|\}$ by excluding from it one of the SRFs $\langle{\bar{\beta}_{q}}|$. We can also define some of the SRFs $\langle\bar{\beta}_{q}^{2}|$ as linear combinations of the SRFs $\langle{\bar{\beta}_{q}}|$: the set $\{\langle\bar{\beta}_{q}^{2}|\}$ should include N1N-1 independent SRFs which can be nonorthogonal. We note that any superposition of $v$ linearly independent SRFs $\beta_{q}(\vec{r})$, q=0,,v1q=0,...\,,v-1, entering Eq. (1) should produce exactly the same results while superpositions involving additional $w$ SRFs $\bar{\beta}_{q}(\vec{r})$}, q=v,,v+wq=v,...\,,v+w, involved in Eq. (22) may affect the convergence. Next, we obtain in the same manner the set $\{\langle\bar{\beta}_{q}^{3}|\}$ including N2N-2 independent SRFs $\langle\bar{\beta}_{q}^{3}|$, etc. More generally, our prescription for SRF selection is as follows: for the first channel, we use a set of $N$ SRFs $\{\langle\bar{\beta}_{q}^{1}|\}$, where q=0,,N1q=0,...,N-1. In the $i$th channel, where i=2,,wi=2,...,w, we remove the SRF $\bar{\beta}_{N-i-w+1}^{i-1}(\vec{r})$ from the (i1)(i-1)th set to form the set of SRFs $\{\langle\bar{\beta}_{q}^{i}|\}$, where q=0,,Niq=0,...,N-i. We note that the set of ket SRFs $\{|\beta_{q^{\prime}}\rangle\}$ is the same in all entrance channels.
@@
$B^{1}_{q}=-\langle\bar{\beta}^{1}_{q}|H-E|\eta^{-}_{{i}}\rangle,\;\qquad$ $\hskip-12.91663ptq=0,...\,,N-1,$ (34c)
@@
[…]